OpenStudy (anonymous):

What are both pairs of integers (x,y) for 4^y-615=x^2. Please show the steps!

OpenStudy (anonymous):

\[ 4^y -615 = x^2 \]This?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Well, consider that \[ 4^y−615=x^2 \\ \sqrt{4^y−615}=\sqrt{x^2} \\ \sqrt{4^y−615}=|x| \\ \pm \sqrt{4^y−615}=x \\ \]So if \((x, y)\) is a pair, then so is \((-x, y)\).

OpenStudy (anonymous):

My next suggestion is to find a prime factorization of \(615\).

OpenStudy (anonymous):

\(615= 3 \cdot 5\cdot 41\)

OpenStudy (anonymous):

You want \(4^y - 3\cdot 5 \cdot 41\) to become a perfect square somehow.

OpenStudy (anonymous):

o, now we can plug in numbers for y? Trial and error?

OpenStudy (anonymous):

Hmmm, well I'm not sure of any mathmatical method for it other than trial and error right now but...

OpenStudy (anonymous):

\[ 4^y−615=x^2 \\ \log_4{(4^y - 615)} = \log_4(x^2) \\ \log_4{(4^y)} / \log_4{(615)}= \log_4(x^2) \\ y\cdot \log_4{(4)} = 2\log_4(x) \cdot \log_4{(615)} \\ y = 2\log_4(x) \cdot \log_4{(615)} \\ \] Another way of looking at the problem

OpenStudy (anonymous):

Hmmm, haven't they told you any methods for problems like these in your class?

OpenStudy (anonymous):

Well, this chapter is on factoring difference of squares and cube, but i cant find the connection

OpenStudy (anonymous):

Also i am a little confuse with the second method because i haven't learn log yet.

OpenStudy (anonymous):

Are you sure it has \(4^y\)? That doesn't make sense then.

OpenStudy (anonymous):

\(4y\) would make more sense.

OpenStudy (anonymous):

it is 4^y

OpenStudy (anonymous):

There is a method is am thinking of that worth trying. Which i learned in class to solve other problems in this chapter. \[4^{y}=2^{2y}=2y^{2}\] \[2y^{2}-x^{2}=615\] \[(2^{y}+x)(2^{y}-x)=615\]

OpenStudy (anonymous):

factors of 615: 1,615 3,123 5,41

OpenStudy (anonymous):

\[2^{y}+x=41\] and \[2^{y}-x=5\], we will be able to find the value for y, right?

OpenStudy (anonymous):

But y won't be an integer

OpenStudy (anonymous):

Hmmm, let my try something... \[ \begin{array}{rcl} 4^y -615 &=& x^2\\ 2^{2y} -615 &=& x^2\\ (2^y)^2 &=& x^2 +615\\ (2^y)^2 -x^2&=& 615\\ (2^y+x)(2^y -x)&=& 615\\ \end{array} \]

OpenStudy (anonymous):

@yociyoci Make sense?

OpenStudy (anonymous):

I thought of this too but if we set the two binomials to be equal to the set of factors of 615 and then solve for y, y won't be an integer

OpenStudy (anonymous):